# Hear What Matters! Text-conditioned Selective Video-to-Audio Generation

### v1

@@ B.1 Data collection
As mentioned in Section 4.1, we acquired 67 clean, mono-source videos through automatic filtering and manual curation. These videos cover 39 unique text labels spanning 8 sound categories in Tab. B2. Fig. B3 summarizes the category-wise statistics. To obtain clean, mono-source audio-video-text samples, we begin with the audio-visual event annotations from UnAV-100 [unav100], a dataset containing timestamped text labels for 100 sound categories across 10,000 videos. First, we identified 907 videos that are common to both UnAV-100 and the VGGSound testset. An automatic filtering step then removed clips annotated with more than one unique sound event. Subsequently, we performed a manual verification process, retaining clips that met three strict criteria: (1) the video contains a single audible sound source with minimal background noise or off-screen sound; (2) the sounding object is clearly visible; and (3) the text annotation precisely matches the sound event. When text annotations from UnAV-100 and VGGSound differed, we manually selected the more appropriate one.
@@ C.1 Baseline models
ReWaS [rewas]
ReWaS leverages a pretrained text-to-audio (TTA) model as its generator for text-conditioned V2A. The model first predicts the audio’s energy curve from the input video and uses this curve as a condition for the TTA model. Since ReWaS natively generates 5-second audio, we adapt it for 8-second videos by splitting each video into two overlapping 5-second segments (0-5s and 3-8s). We generate audio for each segment independently and then construct the final 8-second track by merging the first 4 seconds of the first clip (0-4s) with the last 4 seconds of the second clip (4-8s). We use the official implementation with default parameters [^3] https://github.com/naver-ai/rewas.
VinTAGe [vintage]
VinTAGe is also a text-conditioned V2A model that aims to generate both on-screen and off-screen sounds that are semantically consistent with the text and video. As the model generates 10-second audio, we take the first 8 seconds for evaluation. We use the official code and default parameters for ODE sampling during inference [^4] https://github.com/sakshamsingh1/vintage_aud_gen.
@@
To implement segmentation-based models [hearyourclick, saganet] in our experimental setup, we employ the SoTA video object segmentation model, DEVA [deva], and multimodal-conditioned audio generator model, MMAudio [mmaudio]. Similar to SelVA, this VOS-based pipeline takes a video and a text prompt as condition inputs to improve user controllability. DEVA first predicts a segmentation mask for each frame based on the text prompt. Pixels outside this predicted mask are zeroed out to form a masked video. Therefore, ideally, only the text-related target object is visible. This masked video is subsequently fed into MMAudio, along with the original text prompt, to generate the corresponding audio. We used the official DEVA implementation  [^5] https://github.com/hkchengrex/Tracking-Anything-with-DEVA with its default hyperparameters, which include leveraging SAM [sam] for segmentation, applying semi-online temporal fusion of segmentation hypotheses, and disabling video resizing.
@@ C.2 Metrics
To assess overall audio quality, we adopt three different metrics. Fréchet Audio Distance (FAD) [fad] measures the Fréchet distance between Gaussian distributions fitted to audio embeddings from a reference set and a generated set. Kernel Audio Distance (KAD) [kad], proposed as an unbiased and distribution-free alternative to FAD, also measures this set-wise embedding distance using the Maximum Mean Discrepancy (MMD) with a Gaussian RBF kernel. Inception Score (IS) [inception_score] evaluates both the quality and diversity of generated samples by calculating the KL divergence between the conditional label distribution for individual samples and the marginal distribution across all samples.
For semantic alignment, we report KL divergence, CLAP [clap_laion], and ImageBind [imagebind] scores. The Kullback-Leibler divergence (KL) measures audio semantic similarity using the audio classification distributions of the generated and ground-truth audio. CLAP and IB scores capture global semantic similarity using cosine distance between text-audio and video-audio pairs, respectively.
Finally, to assess audio-video temporal alignment, we report the DeSync score [mmaudio], which is defined as the average predicted offset (in seconds) between the audio and video predicted by a pretrained Synchformer [synchformer].
We use a pretrained PANNs [panns] model to extract audio embeddings for FAD, KAD, IS, and KL, as the model’s features have shown a high correlation with human perception of audio quality [correlation_fad, kad]. All metrics were calculated using open-source toolkits, including av-benchmark  [^6] https://github.com/hkchengrex/av-benchmark and kadtk [^7] https://github.com/YoonjinXD/kadtk.
@@ D.2 VGGSound test set
Tab. D4 summarizes the performance of state-of-the-art text-conditioned V2A models on the VGGSound [vggsound] test set. SelVA achieves results comparable to MMAudio, showing improved semantic alignment but slightly lower temporal alignment. This stems from the nature of VGGSound original test set, which is not curated for selective sound generation and often contains text-irrelevant sound events in videos. Consequently, DeSync may favor holistic generation models (i.e., MMAudio) that reproduce these extraneous sounds over selective generation methods (i.e., SelVA). ReWaS [rewas] and VinTAGe [vintage] underperform in all aspects, particularly in temporal alignment. This is likely because they rely heavily on the text modality: ReWaS leverages a pretrained text-to-audio model, while VinTAGe is trained to generate both on-screen and off-screen sounds based on text descriptions. Additionally, we observe that FAD follows the trend of KAD in Tab. D4 (dataset size: 15k), unlike in Tab. 1 (dataset size: 0.5k). This discrepancy arises because FAD is a biased estimator sensitive to sample size.
@@
ReWaS [rewas] 19.96 1.626   7.66 2.42 0.182 0.1825 1.275
VinTAGe [vintage] 15.96 1.185   8.30 4.91 0.217 0.0486 1.263
MMAudio-S-16k [mmaudio] 7.85 0.338 18.95 1.75 0.235 0.2670 0.492
@@
Table D4: Performance of state-of-the-art models on VGGSound [vggsound] test set. Even though SelVA outperforms those methods on VGG-MonoAudio, SelVA still shows comparable performance on noisy VGGsound test samples.
D.3 More qualitative examples
@@
Fig. 7(c) showcases the semantic-level, cross-modal understanding of SelVA. This intra-class example pairs a target “baby crying” video with a “child singing” video. The task requires the model to semantically ground the text prompt, ignoring the visually present but undesired “child singing” event. Both MMAudio and the VOS baseline fail, generating mumble sounds synchronized with the non-target child on the right. This failure is expected for the VOS baseline, as DEVA [deva] performs object-level segmentation and cannot semantically distinguish between the two subjects based on the text. In contrast, SelVA successfully leverages its text-conditioned vision encoder to generate the correct, synchronized crying sound.
D.4 Limitation of CLAP score
Fig. D8 highlights a limitation of the CLAP score [clap_laion] in capturing semantic text-audio alignment compared to human perception. In general, the VOS baseline was more likely to generate off-screen sounds, an error we attribute to the lack of background pixel information. Such artifacts may not be captured by the CLAP score. We argue that the CLAP encoder trained on noisy audio-text pairs may not penalize the presence of such non-diegetic sounds if they occurred frequently in its training data. However, we found that human annotators are highly sensitive to those artifacts, in that CLAP (0.344) of VOS is comparable to that of SelVA (0.349), but the temporal alignment scores (TA) differ substantially. This result demonstrates that human study is still essential to access V2A generation methods.

### v2

@@ B.1 Data collection
As mentioned in Section 4.1, we acquired 67 clean, mono-source videos through automatic filtering and manual curation. These videos cover 39 unique text labels spanning 8 sound categories in Tab. B2. Fig. B3 summarizes the category-wise statistics. To obtain clean, mono-source audio-video-text samples, we begin with the audio-visual event annotations from UnAV-100 [22], a dataset containing timestamped text labels for 100 sound categories across 10,000 videos. First, we identified 907 videos that are common to both UnAV-100 and the VGGSound testset. An automatic filtering step then removed clips annotated with more than one unique sound event. Subsequently, we performed a manual verification process, retaining clips that met three strict criteria: (1) the video contains a single audible sound source with minimal background noise or off-screen sound; (2) the sounding object is clearly visible; and (3) the text annotation precisely matches the sound event. When text annotations from UnAV-100 and VGGSound differed, we manually selected the more appropriate one.
@@ C.1 Baseline models
ReWaS [32]
ReWaS leverages a pretrained text-to-audio (TTA) model as its generator for text-conditioned V2A. The model first predicts the audio’s energy curve from the input video and uses this curve as a condition for the TTA model. Since ReWaS natively generates 5-second audio, we adapt it for 8-second videos by splitting each video into two overlapping 5-second segments (0-5s and 3-8s). We generate audio for each segment independently and then construct the final 8-second track by merging the first 4 seconds of the first clip (0-4s) with the last 4 seconds of the second clip (4-8s). We use the official implementation [^3] https://github.com/naver-ai/rewas with default parameters.
VinTAGe [39]
VinTAGe is also a text-conditioned V2A model that aims to generate both on-screen and off-screen sounds that are semantically consistent with the text and video. As the model generates 10-second audio, we take the first 8 seconds for evaluation. We use the official code [^4] https://github.com/sakshamsingh1/vintage_aud_gen and default parameters for ODE sampling during inference.
@@
To implement segmentation-based models [45, 60] in our experimental setup, we employ the SoTA video object segmentation model, DEVA [9], and multimodal-conditioned audio generator model, MMAudio [8]. Similar to SelVA, this VOS-based pipeline takes a video and a text prompt as condition inputs to improve user controllability. DEVA first predicts a segmentation mask for each frame based on the text prompt. Pixels outside this predicted mask are zeroed out to form a masked video. Therefore, ideally, only the text-related target object is visible. This masked video is subsequently fed into MMAudio, along with the original text prompt, to generate the corresponding audio. We used the official DEVA implementation  [^5] https://github.com/hkchengrex/Tracking-Anything-with-DEVA with its default hyperparameters, which include leveraging SAM [36] for segmentation, applying semi-online temporal fusion of segmentation hypotheses, and disabling video resizing.
@@ C.2 Metrics
To assess overall audio quality, we adopt three different metrics. Fréchet Audio Distance (FAD) [34] measures the Fréchet distance between Gaussian distributions fitted to audio embeddings from a reference set and a generated set. Kernel Audio Distance (KAD) [11], proposed as an unbiased and distribution-free alternative to FAD, also measures this set-wise embedding distance using the Maximum Mean Discrepancy (MMD) with a Gaussian RBF kernel. Inception Score (IS) [56] evaluates both the quality and diversity of generated samples by calculating the KL divergence between the conditional label distribution for individual samples and the marginal distribution across all samples.
For semantic alignment, we report KL divergence, CLAP [65], and ImageBind [23] scores. The Kullback-Leibler divergence (KL) measures audio semantic similarity using the audio classification distributions of the generated and ground-truth audio. CLAP and IB scores capture global semantic similarity using cosine distance between text-audio and video-audio pairs, respectively.
Finally, to assess audio-video temporal alignment, we report the DeSync score [8], which is defined as the average predicted offset (in seconds) between the audio and video predicted by a pretrained Synchformer [31].
We use a pretrained PANNs [37] model to extract audio embeddings for FAD, KAD, IS, and KL, as the model’s features have shown a high correlation with human perception of audio quality [58, 11]. All metrics were calculated using open-source toolkits, including av-benchmark  [^6] https://github.com/hkchengrex/av-benchmark and kadtk  [^7] https://github.com/YoonjinXD/kadtk.
@@ D.2 VGGSound test set
Tab. D4 summarizes the performance of state-of-the-art text-conditioned V2A models on the VGGSound [5] test set. SelVA achieves results comparable to MMAudio, showing improved semantic alignment but slightly lower temporal alignment. This stems from the nature of VGGSound original test set, which is not curated for selective sound generation and often contains text-irrelevant sound events in videos. Consequently, DeSync may favor holistic generation models (i.e., MMAudio) that reproduce these extraneous sounds over selective generation methods (i.e., SelVA). ReWaS [32] and VinTAGe [39] underperform in all aspects, particularly in temporal alignment. This is likely because they rely heavily on the text modality: ReWaS leverages a pretrained text-to-audio model, while VinTAGe is trained to generate both on-screen and off-screen sounds based on text descriptions. Additionally, we observe that FAD follows the trend of KAD in Tab. D4 (dataset size: 15k), unlike in Tab. 1 (dataset size: 0.5k). This discrepancy arises because FAD is a biased estimator sensitive to sample size.
@@
ReWaS [32] 19.96 1.626   7.66 2.42 0.182 0.1825 1.275
VinTAGe [39] 15.96 1.185   8.30 4.91 0.217 0.0486 1.263
MMAudio-S-16k [8] 7.85 0.338 18.95 1.75 0.235 0.2670 0.492
@@
Table D4: Performance of state-of-the-art models on VGGSound [5] test set. Even though SelVA outperforms those methods on VGG-MonoAudio, SelVA still shows comparable performance on noisy VGGsound test samples.
D.3 Scaling up the model
To analyze scalability, we evaluate larger versions of our audio generator $\mathcal{G}$ in Tab. D5. Following the model configurations of MMAudio [8], the parameter counts for our small, medium, and large variants are 157M, 621M, and 1.03B, respectively. While generating audio at a higher sample rate necessitates more parameters, scaling the model generally improves overall performance. Note that the results of SelVA-S-16k are reported for all other experiments. The limited performance gain of SelVA-L-44k likely stems from the restricted VGGSound; scaling to larger datasets could better exploit its capacity.
Model Audio Quality Semantic Alignment Temporal Alignment
[…]
Inter-class
SelVA-S-16k 51.7 0.676 13.07 1.85 0.292 0.3251 0.721
SelVA-S-44k 52.7 0.561 13.93 1.82 0.305 0.3490 0.739
SelVA-M-44k 53.9 0.548 14.53 1.69 0.315 0.3586 0.695
SelVA-L-44k 52.6 0.595 14.28 1.76 0.305 0.3517 0.691
Intra-class
SelVA-S-16k 37.0 0.492 9.62 1.04 0.280 0.3262 0.639
SelVA-S-44k 38.3 0.340 10.22 1.12 0.297 0.3457 0.695
SelVA-M-44k 39.0 0.360 10.34 1.05 0.295 0.3472 0.646
SelVA-L-44k 38.5 0.346 10.31 1.13 0.294 0.3436 0.659
Table D5: Performance of SelVA with different sizes on VGG-MonoAudio. All methods used text prompts corresponding to the target videos. The best scores are shown in bold, and the second-best scores are underlined.
D.4 More qualitative examples
(a) “Cat Meowing”
(b) “Dog Barking”
Figure D7: Attention visualization for [eos] token over real-world video frame in the last block without (left) / with (right) [SUP] tokens. Each subcaption denotes the corresponding target prompt.
Fig. D7 additionally provides an attention visualization of the [eos] token on a real-world video. By introducing our supplementary token [SUP], our model focuses more effectively on the target sound source while suppressing attention toward the paired object.
@@ D.4 More qualitative examples
Fig. 8(c) showcases the semantic-level, cross-modal understanding of SelVA. This intra-class example pairs a target “baby crying” video with a “child singing” video. The task requires the model to semantically ground the text prompt, ignoring the visually present but undesired “child singing” event. Both MMAudio and the VOS baseline fail, generating mumble sounds synchronized with the non-target child on the right. This failure is expected for the VOS baseline, as DEVA [9] performs object-level segmentation and cannot semantically distinguish between the two subjects based on the text. In contrast, SelVA successfully leverages its text-conditioned vision encoder to generate the correct, synchronized crying sound.
D.5 Limitation of CLAP score
Fig. D9 highlights a limitation of the CLAP score [65] in capturing semantic text-audio alignment compared to human perception. In general, the VOS baseline was more likely to generate off-screen sounds, an error we attribute to the lack of background pixel information. Such artifacts may not be captured by the CLAP score. We argue that the CLAP encoder trained on noisy audio-text pairs may not penalize the presence of such non-diegetic sounds if they occurred frequently in its training data. However, we found that human annotators are highly sensitive to those artifacts, in that CLAP (0.344) of VOS is comparable to that of SelVA (0.349), but the temporal alignment scores (TA) differ substantially. This result demonstrates that human study is still essential to access V2A generation methods.